# The separating Noether number of small groups

### v2

@@ 3.1. Dependence on the base field of $\beta_{\mathrm{sep}}^{K}(G)$
Let $K$ be a subfield of a field $L$, and let $V_{K}$ be a $G$-module. Write $V_{L}$ for the $G$-module $L\otimes_{K}V$ over $L$.
  • (i)
    For any non-negative integer $d$, the $L$-vector space $L[V_{L}]^{G}_{d}$ is spanned by its subset $K[V_{K}]^{G}_{d}$.
  • (ii)
    We have the inequality $\beta_{\mathrm{sep}}(G,V_{K})\leq\beta_{\mathrm{sep}}(G,V_{L})$.
  • (iii)
    We have the inequality $\beta_{\mathrm{sep}}^{K}(G)\leq\beta_{\mathrm{sep}}^{L}(G)$.
Proof.
(i) It is well known, and follows from the fact that the space of solutions over $L$ of a system of homogeneous linear equations with coefficients in $K$ is spanned by the subset of solutions over $K$.
(ii) Take any $v,v^{\prime}\in V_{K}\subseteq V_{L}$ with $G\cdot v\neq G\cdot v^{\prime}$. Then there exists an $f\in L[V_{L}]^{G}_{d}$ with $d\leq\beta_{\mathrm{sep}}(G,V_{L})$ such that $f(v)\neq f(v^{\prime})$. It follows by (i) that there exists an $h\in K[V]^{G}_{d}$ with $h(v)\neq h(v^{\prime})$. This clearly shows the desired inequality.
(iii) is an immediate consequence of (ii). ∎
Remark 3.2.
The inequalities in Lemma 3.1 (ii), (iii) may be strict:
(i) First we mention an example in the modular case. Consider the $\mathrm{S}_{n}$-module $\mathbb{F}_{q}^{n}$ endowed with the natural permutation representation, where $q$ is a prime power and $\mathbb{F}_{q}$ is the finite field with $q$ elements. Separating $\mathrm{S}_{n}$-invariants on $\mathbb{F}_{q}^{n}$ were studied (much before the notion of ’separating invariants’ was formally introduced) in [28], [1], and recently in [35], [20]. In particular, we have $\beta_{\mathrm{sep}}(\mathrm{S}_{3},\mathbb{F}_{2}^{3})=2$ by [35, Lemma 4.3] and $\beta_{\mathrm{sep}}(\mathrm{S}_{3},\mathbb{F}_{4}^{3})=3$ by [20, Corollary 4.9].
(ii) An example where the inequality of Lemma 3.1 (ii) is strict and $\mathrm{char}(K)$ does not divide $|G|$ will be provided in the present paper in Proposition 5.6. In the subsequent paper [21] we shall give examples where the inequality of Lemma 3.1 (iii) is strict and $\mathrm{char}(K)$ does not divide $|G|$.
Lemma 3.1 and Remark 3.2 raise the following question (not answered in the present paper):
Question 3.3.
Can the inequality in Lemma 3.1 (ii) be strict for $K$ infinite?
Proposition 3.4.
Let $G$ be a finite abelian group. Then we have
(6) $\beta_{\mathrm{sep}}^{K}(G)\leq\beta_{\mathrm{sep}}^{\mathbb{C}}(G),\text{ with equality when }K\text{ contains an element }$
whose multiplicative order equals the exponent of $\text{whose multiplicative order equals the exponent of }G.$
Proof.
In [18, Corollary 2.6], a characterization purely in terms of $G$ (and not referring to $K$) is given for $\beta_{\mathrm{sep}}^{K}(G)$, valid when $K$ contains an element whose multiplicative order equals the exponent of $G$ (in fact [18] assumes that the base field is algebraically closed, but the proof of the quoted result works verbatim under the present weaker assumption on $K$). So $\beta_{\mathrm{sep}}^{K}(G)=\beta_{\mathrm{sep}}^{\mathbb{C}}(G)$ when $K$ contains an element of multiplicative order $|G|$. In particular, for any $K$ we have $\beta_{\mathrm{sep}}^{\overline{K}}(G)=\beta_{\mathrm{sep}}^{\mathbb{C}}(G)$, where $\overline{K}$ is the algebraic closure of $K$ (recall the running assumption that $|G|$ is invertible in $K$). Note finally that by Lemma 3.1 (iii) we have $\beta_{\mathrm{sep}}^{K}(G)\leq\beta_{\mathrm{sep}}^{\overline{K}}(G)$. ∎
3.2. Comparison with the dependence on the base field of $\beta^{K}(G)$.
In contrast with Remark 3.2, it is shown in [36, Corollary 4.2] that $\beta^{K}(G)$ depends only on $\mathrm{char}(K)$, so
$\beta^{K}(G)=\begin{cases}\beta^{\mathbb{F}_{p}}(G)\text{ if }\mathrm{char}(K)=p>0;\\ \beta^{\mathbb{C}}(G)\text{ if }\mathrm{char}(K)=0.\end{cases}$
Moreover, by [36, Theorem 4.7],
(7) $\beta^{\mathbb{F}_{p}}(G)\geq\beta^{\mathbb{C}}(G)\text{ for all }p\nmid|G|,\text{ with equality for all but finitely many }p.$
It is interesting to compare this with (6), where the inequality for the corresponding separating Noether numbers is proved in the reverse direction. We mention that no example is known for $p$ and $G$ (where $p\nmid|G|$) for which (7) is a strict inequality. In fact (7) holds with equality for all $p\nmid|G|$ when $G$ is abelian, or when $G$ is one of the groups for which the exact value of their Noether number is given in [15]. For sake of completeness of the picture, we repeat that $\beta^{\mathbb{F}_{p}}(G)=\infty$ if $p$ divides $|G|$ by [43], whereas $\beta_{\mathrm{sep}}^{K}(G)\leq|G|$ for any $K$ (including the case when $\mathrm{char}(K)$ divides $|G|$).
3.3. Direct sum decompositions and multigradings
Assume that we have a $G$-module direct sum decomposition
(8) $V=V_{1}\oplus\cdots\oplus V_{k}.$
View the dual space $V_{i}^{*}$ as the subspace of $V^{*}$ consisting of the linear forms vanishing on $V_{j}$ for $j\neq i$. Then $V^{*}=\bigoplus_{i=1}^{k}V_{i}^{*}$. Choose a basis $x_{i}^{(j)}$ ($i=1,\dots,\dim(V_{j})$) in $V_{i}^{*}$, then our set of variables in $K[V]$ is $\{x_{i}^{(j)}\mid\ j=1,\dots,k;\ i=1,\dots,\dim(V_{j})\}$. Associated to the decomposition (8) there is an $\mathbb{N}_{0}^{k}$-grading on $K[V]$: the multihomogeneous component of multidegree $\alpha=(\alpha_{1},\dots,\alpha_{k})$ in $K[V]$ is spanned by the monomials having degree $\alpha_{j}$ in the variables belonging to $V_{j}^{*}$ for each $j=1,\dots,k$. Clearly, the $G$-action on $K[V]$ preserves this multigrading, and thus $K[V]^{G}$ is spanned by multihomogeneous elements. In the sequel whenever we mention a ‘multigrading’ not explicitly defined, this should refer to the multigrading that comes from a direct sum decomposition as in (8), which should be clear from the context. A typical way by which we shall exploit the multigrading is captured by the following obvious statement:
Lemma 3.5.
We have $\beta_{\mathrm{sep}}^{K}(G)\geq d$ if and only if there exists a $G$-module $V=V_{1}\oplus\cdots\oplus V_{k}$, $v,v^{\prime}\in V$, $h\in K[V]^{G}$ with $h(v)\neq h(v^{\prime})$, such that for any multihomogeneous invariant $f\in K[V]^{G}$ with $\deg(f)<d$ we have $f(v)=f(v^{\prime})$.
Note that the above multigrading on $K[V]$ is induced by the following action of the torus $T:=K^{\times}\times\cdots\times K^{\times}$ ($k$ direct factors) on $V$: for $\lambda=(\lambda_{1},\dots,\lambda_{k})\in T$ and $v=(v_{1},\dots,v_{k})\in V$ set $\lambda\cdot v:=(\lambda_{1}v_{1},\dots,\lambda_{k}v_{k})$. This action commutes with the $G$-action on $V$. Consequently, for $v,v^{\prime}\in V$ we have that
Later in the text when we say that ’after rescaling, we may assume that $v$ has some special form’, then we mean that we replace $v$ by an appropriate element in its $T$-orbit, and we refer to (9).
3.4. Reduction to multiplicity-free representations
We record a consequence of the results of [23] and Lemma 3.1:
Lemma 3.6.
Let $V_{1},\dots,V_{q}$ be a complete irredundant list of representatives of the isomorphism classes of simple $G$-modules. Assume that
$|K|>(\max\{\dim(V_{j})\mid j=1,\dots,q\}-1)|G|.$
Then we have the equality
$\beta_{\mathrm{sep}}^{K}(G)=\beta_{\mathrm{sep}}(G,V_{1}\oplus\cdots\oplus V_{q}).$
Proof.
An arbitrary $G$-module $V$ is isomorphic to $V_{1}^{m_{1}}\oplus\cdots\oplus V_{q}^{m_{q}}$, where $W^{m}$ stands for the direct sum $W\oplus\cdots\oplus W$ of $m$ copies of $W$. By [23, Corollary 2.6] we have $\beta_{\mathrm{sep}}(G,V)\leq\beta_{\mathrm{sep}}(G,V_{1}^{\dim(V_{1})}\oplus\cdots\oplus V_{q}^{\dim(V_{q})})$, and by [23, Theorem 3.4 (ii)] and [23, Proposition 3.3 (ii)] we have $\beta_{\mathrm{sep}}(G,V_{1}^{\dim(V_{1})}\oplus\cdots\oplus V_{q}^{\dim(V_{q})})=\beta_{\mathrm{sep}}(G,V_{1}\oplus\cdots\oplus V_{q})$. ∎
Therefore to compute $\beta_{\mathrm{sep}}^{K}(G)$ it is sufficient to deal with multiplicity-free representations of $G$ (unless possibly if $K$ is a finite field with too few elements).
3.5. Helly dimension
Lemma 3.6 can be improved for groups whose Helly dimension is ’small’. The Helly dimension $\kappa(G)$ is the minimal positive integer $k$ having the following property: any set of cosets in $G$ with empty intersection contains a subset of at most $k$ cosets with empty intersection (see [17, Section 4]).
Lemma 3.7.
Under the assumptions and notation of Lemma 3.6, there exists a $k\leq\kappa(G)$ and $1\leq j_{1}<\dots<j_{k}\leq q$ with $\beta_{\mathrm{sep}}^{K}(G)=\beta_{\mathrm{sep}}(G,V_{j_{1}}\oplus\cdots\oplus V_{j_{k}})$.
Proof.
This follows from Lemma 3.6 by the same argument as the proof of [22, Lemma 4.1] (formulated in a bit more general setup). ∎
In view of Lemma 3.7, it is helpful to find upper bounds on the Helly dimension. By [17, Lemma 4.2] we have the inequality
(10) $\kappa(G)\leq\mu(G)+1,$
where $\mu(G)$ stands for the maximal size of an intersection independent set of subgroups of $G$. Recall that a set of subgroups of $G$ is intersection independent if none contains the intersection of the others.
Lemma 3.8.
If a set $S$ of intersection independent subgroups in a group contains a cyclic subgroup of prime power order, then $|S|\leq 2$.
Proof.
Let $S$ be an intersection independent set of subgroups of a group $G$ with $|S|\geq 2$ and $H\in S$ with $|H|$ a prime power. Take $J\in S$ with $J\cap H$ of minimal possible order. Since the lattice of subgroups of $H$ is a chain, for any $L\in S$ we have $L\cap H\supseteq H\cap J$, hence in particular $L\supseteq H\cap J$. Since $S$ is intersection independent, it follows that $L\in\{H,J\}$. As $L\in S$ was arbitrary, it means that $S=\{H,J\}$. ∎
3.6. Notational convention.
Denote by $\widehat{G}$ the group of characters of $G$ (i.e. the group of homomorphisms $G\to K^{\times}$, written multiplicatively). For $\chi\in\widehat{G}$ denote by $U_{\chi}$ the $1$-dimensional vector space $K$ endowed with the representation $\chi$; that is, $g\cdot\lambda=\chi(g)\lambda$ for $g\in G$ and $\lambda\in K$. In several cases we shall deal with a $G$-module
where the $W_{i}$ are pairwise non-isomorphic irreducible $G$-modules of dimension at least $2$, the $\chi,\dots,\chi_{m}$ are distinct characters of $G$. We shall denote by $x_{1},x_{2},\dots$ the coordinate functions on $W_{1}$, by $y_{1},y_{2},\dots$ the coordinate functions on $W_{2}$, $z_{1},z_{2},\dots$ the coordinate functions on $W_{3}$, etc., and denote by $t_{\chi}$ the coordinate function on $U_{\chi}$. According to the above conventions, we have $g\cdot t_{\chi}=\chi(g^{-1})t_{\chi}$. Associated to the above direct sum decomposition of $V$ is an $\mathbb{N}_{0}^{l+m}$-grading on $K[V]$ explained at the beginning of Section 3. The phrase ‘multihomogeneous’ will refer to this multigrading.
Usually we shall specify a concrete representation of $G$ (in other words, a $G$-module) by giving a group homomorphism $\psi:G\to\mathrm{GL}_{n}(K)$. The underlying vector space is then the space of column vectors $V=K^{n}$, on which $g\in G$ operates via multiplication by the matrix $\psi(g)$. The action of $G$ on the variables $x_{1},\dots,x_{n}$ (which form a basis in the dual space $V^{*}$ of $V$) is given explicitly by the following formula:
$g\cdot x_{j}=\sum_{i=1}^{n}\psi(g^{-1})_{ji}x_{i}.$
We shall write $A^{T}$ for the transpose of the matrix (e.g. row vector) $A$.
3.7. Relative invariants and the Hilbert ideal
Given a $G$-module $V$, denote by $K[V]^{G}_{+}$ the sum of the positive degree homogeneous components of $K[V]^{G}$. A homogeneous invariant $f\in K[V]^{G}$ is said to be indecomposable if $f\notin(K[V]^{G}_{+})^{2}$. Clearly, $K[V]^{G}$ is generated by the positive degree homogeneous indecomposable invariants. The so-called Hilbert ideal $\mathcal{H}(G,V)$ is the ideal in the polynomial algebra $K[V]$ generated by the positive degree homogeneous $G$-invariants. The Hilbert ideal plays an essential role in studying generators of the algebra of invariants, because a positive degree homogeneous invariant $f\in K[V]^{G}$ is indecomposable if and only if $f$ is not contained in the ideal $\mathcal{H}(G,V)K[V]_{+}$ of $K[V]$. More precisely, a set of homogeneous invariants form a minimal generating set of $K[V]^{G}$ if and only if it gives a vector space basis in a vector space direct complement of $(K[V]^{G}_{+})^{2}$ in $K[V]^{G}_{+}$ (the sum of the homogeneous components of $K[V]^{G}$ of positive degree).
An element $f\in K[V]$ is a relative invariant of weight $\chi\in\widehat{G}$ if $g\cdot f=\chi(g^{-1})f$ for all $g\in G$ (so $f(gv)=\chi(g)f(v)$ for all $g\in G$ and $v\in V$). We set
$K[V]^{G,\chi}:=\{f\in K[V]\mid f\text{ is a relative invariant of weight }\chi\}.$
For example, if $V$ is as in (11), then the variable $t_{\chi}$ is a relative invariant of weight $\chi$. Moreover, for a sequence $\chi^{(1)},\dots,\chi^{(k)}$ of elements of $\widehat{G}$, $t_{\chi^{(1)}}\cdots t_{\chi^{(k)}}$ is an indecomposable $G$-invariant if and only if $\chi^{(1)}\cdots\chi^{(k)}=1\in\widehat{G}$ and there is no $\ell<k$, $1\leq i_{1}<\cdots<i_{\ell}\leq k$ with $\chi^{(i_{1})}\cdots\chi^{(i_{\ell})}=1\in\widehat{G}$ (we say in this case that $\chi^{(1)},\dots,\chi^{(k)}$ is an irreducible product-one sequence over $\widehat{G}$, and we call $k$ its length). We say that $\chi^{(1)},\dots,\chi^{(k)}$ is a product-one free sequence over $\widehat{G}$ if there is no $1\leq i_{1}<\dots<i_{\ell}\leq k$ with $\chi^{(i_{1})}\cdots\chi^{(i_{\ell})}=1\in\widehat{G}$. The Davenport constant $\mathsf{D}(\widehat{G})$ of the finite abelian group $G$ is defined as the maximal length of an irreducible product-one sequence over $G$. Clearly, the maximal possible length of a product-one free sequence is $\mathsf{D}(\widehat{G})-1$. For more information about the Davenport constant and its relevance for the invariant theory of abelian groups see for example [14].
Lemma 3.9.
Let $V=W\oplus U$ be a $G$-module as in (11). Let $A$ be a homogeneous generating system of $K[W]^{G}$, and for $\chi\in\{\chi_{1},\dots,\chi_{m}\}$ let $A_{\chi}\subset K[W]^{G,\chi}$ be a set of homogeneous relative invariants of weight $\chi$ that span a vector space direct complement in $K[W]^{G,\chi}$ of $\mathcal{H}(G,W)\cap K[W]^{G,\chi}=K[W]^{G}_{+}K[W]^{G,\chi}$. Set
$B:=\{t_{\chi^{(1)}}\cdots t_{\chi^{(k)}}\mid\chi^{(1)},\dots\chi^{(k)}\text{ is an irreducible product-one sequence over }\widehat{G}\},$
and
$C:=\{ht_{\chi^{(1)}}\cdots t_{\chi^{(k)}}\mid\chi^{(1)},\dots\chi^{(k)}\text{ is a product-one free sequence over }\widehat{G},$
@@ 3.7. Relative invariants and the Hilbert ideal
As we pointed out above, the algebra $K[V]^{G}$ is generated by indecomposable multihomogeneous invariants. A multihomogeneous invariant $f$ belongs to $K[W]^{G}$ or $K[U]^{G}$ or is of the form $ht$, where $t$ is a monomial in $K[U]$ of positive degree, and $h$ is a multihomogeneous element of positive degree in $K[W]$. In the first two cases $f$ belongs to the $K$-subalgebra of $K[V]$ generated by $A\cup B$. In the third case $t=t_{\chi^{(1)}}\cdots t_{\chi^{(k)}}$, where $\chi^{(1)},\dots\chi^{(k)}$ is a product-one free sequence over $\widehat{G}$. Then $t$ is a relative $G$-invariant of some weight $\chi$, and hence necessarily, $h\in K[V]^{G,\chi^{-1}}$. There exists a linear combination $a$ of elements from $A_{\chi^{-1}}$ with $h-a\in K[W]^{G}_{+}K[W]^{G,\chi^{-1}}$. Then $at$ belongs to the $K$-subspace of $K[V]$ generated by $C$, and $f-at=(h-a)t\in(K[V]^{G}_{+})^{2}$. So we showed that any indecomposable multihomogenous element of $K[V]^{G}$ can be reduced modulo $(K[V]^{G}_{+})^{2}$ to an element in the $K$-subalgebra of $K[V]$ generated by $A\cup B\cup C$. This clearly implies our statement. ∎
Given a set $S\subset K[V]$ of polynomials write
$\mathcal{V}(S):=\{v\in V\mid\forall f\in S\colon f(v)=0\}$
for the common zero locus in $V$ of the elements of $S$. Next we restate [9, Lemma 4.3] in a sharp form:
@@
We have the equality
$\mathcal{V}(K[V]^{G,\chi})=\{v\in V\mid\mathrm{Stab}_{G}(v)\nsubseteq\ker(\chi)\}.$
@@
therefore $\chi(g)\neq 1$ implies $f(v)=0$. This shows the inclusion $\{v\in V\mid\mathrm{Stab}_{G}(v)\nsubseteq\ker(\chi)\}\subseteq\mathcal{V}(K[V]^{G,\chi}$. The reverse inclusion is stated and proved in [9, Lemma 4.3]. ∎
3.8. The use of automorphisms
The natural action of the automorphism group of $G$ on the isomorphism classes of representations of $G$ helps to save some computations later, thanks to the following statements:
@@ 3.8. The use of automorphisms
Let $(V,\rho)$ be a $G$-module and $\alpha$ an automorphism of the group $G$.
@@
    For any $\chi\in\widehat{G}$ we have
    $K[(V,\rho)]^{G,\chi}=K[(V,\rho\circ\alpha)]^{G,\chi\circ\alpha}.$
    In particular,
    $K[(V,\rho)]^{G}=K[(V,\rho\circ\alpha)]^{G}.$
@@
    We have
    $\beta(G,(V,\rho))=\beta(G,(V,\rho\circ\alpha))\text{ and }\beta_{\mathrm{sep}}(G,(V,\rho))=\beta_{\mathrm{sep}}(G,(V,\rho\circ\alpha)).$
@@
    Assume that $(V,\rho)\cong(V,\rho\circ\alpha)$ as $G$-modules, so there exists a $K$-vector space isomorphism $T:V\to V$ with $T\circ\rho(g)=(\rho\circ\alpha)(g)\circ T$ for all $g\in G$. Then the map $f\mapsto f\circ T$ induces a $K$-vector space isomorphism
    $K[(V,\rho)]^{G,\chi}\cong K[[(V,\rho)]^{G,\chi\circ\alpha}.$

### v3

@@ 2.2. Relative invariants
Let $V=W\oplus U$ be a $KG$-module as in (2). Let $A$ be a homogeneous generating system of $K[W]^{G}$, and for $\chi\in\{\chi_{1},\dots,\chi_{m}\}$ let $A_{\chi}\subset K[W]^{G,\chi}$ be a set of homogeneous relative invariants of weight $\chi$ that span a vector space direct complement in $K[W]^{G,\chi}$ of $\mathcal{H}(G,W)\cap K[W]^{G,\chi}=K[W]^{G}_{+}K[W]^{G,\chi}$. Set
$B:=\{t_{\chi^{(1)}}\cdots t_{\chi^{(k)}}\mid\chi^{(1)},\dots,\chi^{(k)}\text{ is an irreducible product-one sequence over }\widehat{G}\}$,
$C:=\{ht_{\chi^{(1)}}\cdots t_{\chi^{(k)}}\mid\chi^{(1)},\dots,\chi^{(k)}\text{ is a product-one free sequence over }\widehat{G}$,
@@
As we pointed out above, the algebra $K[V]^{G}$ is generated by indecomposable multihomogeneous invariants. A multihomogeneous invariant $f$ belongs to $K[W]^{G}$ or $K[U]^{G}$ or is of the form $ht$, where $t$ is a monomial in $K[U]$ of positive degree, and $h$ is a multihomogeneous element of positive degree in $K[W]$. In the first two cases $f$ belongs to the $K$-subalgebra of $K[V]$ generated by $A\cup B$. In the third case $t=t_{\chi^{(1)}}\cdots t_{\chi^{(k)}}$, where $\chi^{(1)},\dots,\chi^{(k)}$ is a product-one free sequence over $\widehat{G}$. Then $t$ is a relative $G$-invariant of some weight $\chi$, and hence necessarily, $h\in K[V]^{G,\chi^{-1}}$. There exists a linear combination $a$ of elements from $A_{\chi^{-1}}$ with $h-a\in K[W]^{G}_{+}K[W]^{G,\chi^{-1}}$. Then $at$ belongs to the $K$-subspace of $K[V]$ generated by $C$, and $f-at=(h-a)t\in(K[V]^{G}_{+})^{2}$. So we showed that any indecomposable multihomogenous element of $K[V]^{G}$ can be reduced modulo $(K[V]^{G}_{+})^{2}$ to an element in the $K$-subalgebra of $K[V]$ generated by $A\cup B\cup C$. This clearly implies our statement. ∎
Given a set $S\subset K[V]$ of polynomials write $\mathcal{V}(S):=\{v\in V\mid\forall f\in S\colon f(v)=0\}$ for the common zero locus in $V$ of the elements of $S$. Next we restate [5, Lemma 4.3] in a sharp form:
@@
We have the equality $\mathcal{V}(K[V]^{G,\chi})=\{v\in V\mid\mathrm{Stab}_{G}(v)\nsubseteq\ker(\chi)\}$.
@@
therefore $\chi(g)\neq 1$ implies $f(v)=0$. This shows the inclusion $\{v\in V\mid\mathrm{Stab}_{G}(v)\nsubseteq\ker(\chi)\}\subseteq\mathcal{V}(K[V]^{G,\chi})$. The reverse inclusion is stated and proved in [5, Lemma 4.3]. ∎
2.3. Some reductions
Both the Noether number and the separating Noether number are monotone for subgroups as well; that is, for any subgroup $H$ of a finite group $G$ we have
[…]
@@ 2.3. Some reductions
[15, Lemma 2.4 (iii)] For any subfield $K$ of a field $L$ we have the inequality $\beta_{\mathrm{sep}}^{K}(G)\leq\beta_{\mathrm{sep}}^{L}(G)$.
A consequence of the results of [16] is that if $K$ is a field with sufficiently many elements, one can compute compute $\beta_{\mathrm{sep}}^{K}(G)$ by dealing only with multiplicity-free representations.
Lemma 2.4.
[15, Lemma 6.1] Let $V_{1},\dots,V_{q}$ be a complete irredundant list of representatives of the isomorphism classes of simple $KG$-modules. Assume that
$|K|>(\max\{\dim(V_{j})\mid j=1,\dots,q\}-1)|G|.$
Then we have the equality $\beta_{\mathrm{sep}}^{K}(G)=\beta_{\mathrm{sep}}(G,V_{1}\oplus\cdots\oplus V_{q})$.
Given a group homomorphism $\rho$ from $G$ into the group $\mathrm{GL}(V)$ of invertible linear transformations of $V$, we write $(V,\rho)$ for the $KG$-module $V$ with $g\cdot v:=\rho(g)(v)$ for $g\in G$, $v\in V$. We make use of the natural action of the automorphism group of $G$ on the isomorphism classes of representations of $G$ through the following statement:
Lemma 2.5.
[15, Lemma 6.4] Let $(V,\rho)$ be a $KG$-module and $\alpha$ an automorphism of the group $G$.
@@
    For any $\chi\in\widehat{G}$ we have $K[(V,\rho)]^{G,\chi}=K[(V,\rho\circ\alpha)]^{G,\chi\circ\alpha}$.
@@
    We have $\beta(G,(V,\rho))=\beta(G,(V,\rho\circ\alpha))\text{ and }\beta_{\mathrm{sep}}(G,(V,\rho))=\beta_{\mathrm{sep}}(G,(V,\rho\circ\alpha))$.
@@
    Assume that $(V,\rho)\cong(V,\rho\circ\alpha)$ as $KG$-modules, so there exists a $K$-vector space isomorphism $T:V\to V$ with $T\circ\rho(g)=(\rho\circ\alpha)(g)\circ T$ for all $g\in G$. Then the map $f\mapsto f\circ T$ induces a $K$-vector space isomorphism
    $K[(V,\rho)]^{G,\chi}\cong K[(V,\rho)]^{G,\chi\circ\alpha}.$
3. Groups $G$ with $\beta_{\mathrm{sep}}^{K}(G)=\beta^{K}(G)$
In this subsection
Assume that $K$ contains an element $\xi$ of multiplicative order $6$. So $\omega:=\xi^{2}$ has multiplicative order $3$. Consider the following irreducible $2$-dimensional representations of $G$:
Proposition 3.1.
We have $\beta_{\mathrm{sep}}(G,W_{1}\oplus W_{2})\geq 6$.